OpenStudy (anonymous):

Match the sequence (term) with the correct type of sequence (definition). Match Term Definition 2, 4, 6, 8, ... A) Geometric, common ratio is 2 28, 24, 20, 16, ... B) Geometric, common ratio is 0.5 2, 4, 8, 16, ... C) Arithmetic, common difference is -4 48, 24, 12, 6, ... D) Arithmetic, common difference is 2

OpenStudy (anonymous):

2, 4, 6, 8 is A or D which one?

OpenStudy (amoodarya):

2, 4, 6, 8, ... is arithmetic because every time add +2

OpenStudy (anonymous):

Okay what about the others?

OpenStudy (amoodarya):

2, 4, 8, 16, ... is geometric because every time multiplied by 2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

after the last ones i have another question

OpenStudy (amoodarya):

28, 24, 20, 16, ... is arithmetic because every time add by -4

OpenStudy (amoodarya):

so the answer is ?

OpenStudy (anonymous):

E.J. has shown that a function, f(x), grows by 12% for every unit in the domain. What does this prove? The function f(x) is an arithmetic sequence. The function f(x) is a geometric sequence. The function f(x) is not a sequence. This does not prove anything.

OpenStudy (amoodarya):

48, 24, 12, 6, ... is geometric ,ratio is 1/2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

E.J. has shown that a function, f(x), grows by 12% for every unit in the domain. What does this prove? The function f(x) is an arithmetic sequence. The function f(x) is a geometric sequence. The function f(x) is not a sequence. This does not prove anything.

OpenStudy (anonymous):

this is the last one

OpenStudy (amoodarya):

The function f(x) is an arithmetic sequence when a(n)=a1+(n-1)d it is linear relation (like a line)

OpenStudy (anonymous):

Yes, i think

OpenStudy (amoodarya):

The function f(x) is an geometric sequence when a(n)= a1* q^(n-1)

OpenStudy (anonymous):

So the answer is B

OpenStudy (amoodarya):

I think no one of them is correct , may be you type mistake!

OpenStudy (anonymous):

I copyed it and pasted

OpenStudy (anonymous):

copied

OpenStudy (amoodarya):

2, 4, 6, 8, ... A) Geometric, common ratio is 2 arithmetic +2 28, 24, 20, 16, ... B) Geometric, common ratio is 0.5 arithmetic -4 2, 4, 8, 16, ... C) Arithmetic, common difference is -4 geometric +2 48, 24, 12, 6, ... D) Arithmetic, common difference is 2 geometric 1/2

OpenStudy (anonymous):

E.J. has shown that a function, f(x), grows by 12% for every unit in the domain. What does this prove? The function f(x) is an arithmetic sequence. The function f(x) is a geometric sequence. The function f(x) is not a sequence. This does not prove anything. (I need help with this now sir)

OpenStudy (amoodarya):

a a+12/100 a = (1.12)a (1.12)a +12/100 (1.12)a=(1.12)a (1+.12) = (1.12)^2 a (1.12)^2 a +12/100 (1.12)^2 a= (1.12)^3 a ... for n times (1.12)^n a